# Unmasking Deepfakes: Leveraging Augmentations and Features Variability for Deepfake Speech Detection

### v1

@@ 3 Results and Insights
This section presents the experimental results of the proposed methods, beginning with an overview of the experimental setup and evaluation metrics. We then evaluate the architecture’s performance and training process while experimenting with various augmentation techniques, focusing on examining the role of MaskedSpec and MaskedFeature augmentations in improving model robustness and generalization across diverse conditions. Next, we examine the role of pretraining the feature extractor in improving representation quality. Finally, we performed a fusion of the top-performing models, each trained with distinct augmentation strategies and pretraining configurations.Models trained with MaskedSpec using the Bands mask and MaskedFeature with the Gauss mask, alongside different pretraining weights for the feature extractor, were fused to achieve an EER of 3.39%. This result highlights the value of integrating our independent augmentation strategies that capture diverse perspectives on the data during training, thereby enhancing the model’s robustness and overall performance.
@@ 3.2 Model Architecture and Features
The results in Table 1 demonstrate the clear advantages of trainable features over traditional spectrogram-based representations. Models trained with spectrogram features exhibit the weakest performance, with a Dev EER of 47.86% and Eval EER of 38.72%. This indicates that spectrograms, without additional learned enhancements, struggle to effectively separate genuine from spoofed audio in this task. Introducing Wav2Vec2 as the feature extractor significantly reduces the EER, even without pretraining, achieving a Dev EER of 38.24% and Eval EER of 40.64%. Despite this improvement, the lack of domain-specific pretraining likely limits its generalization capabilities.
Significant improvement is observed when the model is trained in two distinct phases. Initially, Wav2Vec2 undergoes self-supervised pertaining, as detailed in [22], on the ASVSpoof5 dataset, allowing it to learn meaningful representations of audio data without relying on labels. Subsequently, end-to-end training is performed on the hybrid model using labeled data, fine-tuning both the feature extractor and the classification head for the specific task. This dual-phase training approach achieves a notable reduction in error rates, resulting in a Dev EER of 20.82% and an Eval EER of 26.36%, showcasing the combined benefits of unsupervised pretraining and supervised fine-tuning.
@@ 3.3 MaskedSpec and MaskedFeature Experiments
Squares 2.4972 6.73
Bands 1.4397 8.78
Singles 2.0733 13.13
Gauss 1.3802 5.22
@@
3.4 Pretraining Feature Extractor
Feature Extractor Pretrainig Compressions $MS_{B}$ $MF_{G}$ Dev EER (%) Eval EER (%)
@@ 3.4 Pretraining Feature Extractor
ASVSpoof5+ 5.33 5.11
5.96 5.25
Table 5: Evaluation of model performance across various pretrained feature extractors and augmentation strategies. The overall best results are highlighted in bold, with the top-performing results for the ASVspoof 5 Track 1 Closed Condition marked in light gray
We conducted end-to-end training of models initialized with various pretrained weights, using the top-performing configurations for MaskedSpec and MaskedFeature augmentations: Bands and Gauss, respectively. These augmentation configurations are referred to as $MS_{B}$ and $MF_{G}$ in the following sections. The experimental results are summarized in Table 5. To establish a baseline for comparison with subsequent experiments, we first evaluate the model’s performance without augmentation or pertaining the feature extractor. This provides a reference point for assessing the impact of various enhancements.
We investigate the effects of pretraining the feature extractor using XLS-R 53 and ASVSpoof5, both with and without compression augmentations. XLS-R 53 utilizes publicly available weights trained on a diverse dataset, including Multilingual LibriSpeech, CommonVoice, and Babel, comprising over 50k hours of audio. We pretrained a feature extraction model using the ASVSpoof5 dataset, which contains bona fide and spoofed audio in English only and is significantly smaller in scale compared to the datasets used for XLS-R 53. As expected, XLS-R 53 pretrining substantially outperforms ASVSpoof5 pretrining due to its larger and more diverse data. Furthermore, the inclusion of compression augmentations consistently improves performance across all configurations.
Furthermore, we trained the model using the $MS_{B}$ and $MF_{G}$ augmentations. We experimented with additional pretrained weights, including XLS-R 128 [24], which was trained on 436k hours of unlabeled speech data across 128 languages, leveraging datasets such as VoxPopuli, MLS, CommonVoice, BABEL, and VoxLingua107. Another feature extractor. Large-960h, was trained on 960 hours of English-only LibriSpeech data. Additionally, we pretrained with ASVSpoof5 extended with codec augmentations, referred to as ASVSpoof5+.
The results underscore the substantial impact of the pretraining and the augmentation on model performance. Using XLS-R 53 weights, training with $MS_{B}$ and $MF_{G}$ augmentations yielded performance improvements of 26.34% and 15.12%, respectively, compared to models trained with compression augmentations only. Even greater improvements were observed with ASVSpoof5 pretraining, where $MS_{B}$ and $MF_{G}$ augmentations enhanced performance by 38.27% and 30.91%, respectively. Notably, incorporating compression augmentations during pretraining further improved performance, achieving significant gains of 75.08% and 74.4% with $MS_{B}$ and $MF_{G}$, respectively.
Incorporating compression augmentations into pretraining demonstrated a notable impact, yielding approximately a 60% improvement in model performance. This enhancement was observed when comparing models pretrained with ASVSpoof5 to those pretrained with ASVSpoof5+ using either $MS_{B}$ or $MF_{G}$ augmentations.
3.5 Fusions
ASVSpoof5+ XLS-R 53 XLS-R 128 Eval EER (%)
$MS_{B}$ $MF_{G}$ $MS_{B}$ $MF_{G}$ $MS_{B}$ $MF_{G}$
4.37
3.61
3.39
Table 6: Fusions results.
The results presented in Table 6 demonstrate the performance of various fusions of the single models presented in Table 5. The best evaluation EER of 3.39% is achieved by combining models where the feature extractors were pretrained on XLS-R 53, XLS-R 128, and ASVSpoof5+, utilizing both $MS_{B}$ and $MF_{G}$ augmentations. Under ASVSpoof5 closed-conditions, the fusion achieved an EER of 4.37% using $MS_{B}$ and $MF_{G}$ augmentations with feature extractor pretrained on ASVSpoof5+, establishing SoTA results for this benchmark.
4 Analysis of Model Performance Across Deepfake Attacks
Our assessment of model performance relied on insights provided by the competition organizers. These insights included trends in EER across different codecs and spoofing attacks, enabling us to gauge model efficacy without direct access to specific labels. Our observations reveal variation in detection difficulty across attack types, with certain spoofing techniques consistently associated with lower EERs, indicating they are more readily identified by most models.
@@ 4 Analysis of Model Performance Across Deepfake Attacks
XLS-R 128, $MS_{B}$ 1.40 3.86 2.23 2.04 0.93 2.10 3.61 3.87 1.65 1.61 4.79 10.78 0.99 6.95 9.20 3.16
XLS-R 128, $MF_{G}$ 2.37 6.66 3.39 3.59 1.89 2.82 3.79 6.22 2.32 2.77 9.44 12.57 1.78 12.98 12.43 7.14
XLS-R 53, $MS_{B}$ 1.62 2.68 1.32 1.33 1.56 2.34 3.53 6.03 1.42 1.41 1.76 17.32 1.42 3.76 5.10 1.47
XLS-R 53, $MF_{G}$ 2.00 4.53 2.01 1.75 2.22 2.58 5.05 8.71 1.63 1.94 2.09 16.30 1.74 5.52 6.60 1.56
ASVSpoof5+, $MS_{B}$ 1.70 0.72 5.61 5.01 2.03 4.68 1.64 7.80 3.98 1.29 1.10 19.48 0.84 1.27 4.40 1.66
ASVSpoof5+, $MF_{G}$ 2.19 0.85 5.69 5.25 3.30 5.48 3.37 7.67 4.44 1.84 1.00 22.79 1.04 1.46 4.36 1.52
ASVSpoof5+ fusion 1.19 0.57 3.98 3.72 1.87 3.60 1.91 6.61 3.05 1.00 0.74 19.53 0.74 0.90 3.24 1.02
fusion 1.01 1.82 1.36 1.27 1.23 1.60 2.27 3.50 1.10 1.20 1.79 8.77 0.92 2.53 4.86 1.25
Table 7: Pooled EER performance analysis for spoof attacks. The lowest EER for each spoof attack among single models is highlighted in bold, while the second-lowest is marked in light gray.
The deepfake attacks, denoted A17–A32, presented a range of detection challenges. Attacks A17, A19, A20, A21, A25, A26, and A29 generally produced low EERs across models, suggesting either a relative ease of detection or high compatibility of augmentation methods with these attack types. In contrast, attacks A28, A30, and A31 posed significant detection challenges, as indicated by higher EERs across the majority of tested models. In particular, attack A28 yielded notably elevated EERs, highlighting it as one of the most challenging spoofing types. As presented in Table 7, models utilizing XLS-R 53 pretrained feature extractor, trained with $MS_{B}$, demonstrated robust performance across a broad spectrum of spoofing methods. ASVSpoof5+ achieved lowest EERs on A30 and A31. For A28, identified as the most difficult spoofing type, the model trained with XLS-R 128 pretrained feature extractor and $MS_{B}$ achieved the lowest EER, closely followed by the XLS-R 128 trained with $MF_{G}$. These findings underscore the benefits of augmentations on both audio spectrogram and feature levels, as well as the importance of leveraging pretrained feature extractor weights for improved deepfake speech detection.
When analyzing the fusion of all single models, it is evident that this approach significantly improves performance for attacks such as A17, A20, A22, A24-26, A28, and A32, achieving the lowest EER compared to any single model. For other spoofing attacks, the performance of the fusion model remains comparable to the lowest EER achieved by a single model. Interestingly, the ASVSpoof5+ fusion outperforms the full fusion model on A18, A23, A26-27, A29-32, but demonstrates considerably poorer performance on A28.
Model None C1 C10 C11 C2 C3 C4 C5 C6 C7 C8 C9
XLS-R 128, $MS_{B}$ 1.73 3.10 5.57 2.11 2.89 4.29 10.88 1.96 2.05 13.83 5.90 3.76
XLS-R 128, $MF_{G}$ 3.22 6.02 9.94 3.88 7.07 7.31 12.77 3.31 4.09 15.38 7.58 7.75
XLS-R 53, $MS_{B}$ 1.52 3.47 5.88 3.23 2.97 4.39 7.92 1.69 2.65 9.77 6.37 4.38
XLS-R 53, $MF_{G}$ 1.94 3.61 6.80 3.67 3.52 5.05 9.97 2.00 2.62 12.36 6.72 4.97
ASVSpoof5+, $MS_{B}$ 3.60 4.73 6.49 3.95 4.74 5.69 6.41 4.38 4.67 7.28 6.94 5.77
ASVSpoof5+, $MF_{G}$ 3.88 5.47 7.53 4.62 5.44 7.66 7.30 4.25 4.68 7.81 7.51 6.36
ASVSpoof5+ fusion 3.03 4.26 6.15 3.64 4.26 5.69 5.92 3.59 3.90 6.54 6.23 5.28
fusion 0.59 1.62 3.77 1.21 1.59 2.39 7.53 0.74 0.89 9.55 3.70 2.38
@@
Table 8: Pooled EER performance analysis for codecs. The lowest EER for each codec among single models is highlighted in bold, while the second-lowest is marked in light gray.

### v2

@@ 3 Results and Insights
This section presents a detailed evaluation of the proposed deepfake detection framework, focusing on the contribution of individual components and the impact of the proposed augmentation strategies. We begin by assessing the effect of various masking configurations at both the spectrogram and feature levels. We then evaluate the role of compression and RawBoost augmentations under different training regimes. Fusion analyses and ablation studies are included to further understand the individual and joint effects of the system components. Performance is reported primarily in terms of Equal Error Rate (EER), additional results, including evaluations using minimum Detection Cost Function (minDCF), are provided in the Appendix for completeness.
@@ 3.2 Model Architecture and Features
Trainable SSL features substantially outperform traditional spectrogram inputs, and self-supervised pretraining provides a significant boost in both development and evaluation EER.
@@
The results in Table 1 demonstrate the clear advantages of trainable features over traditional spectrogram-based representations. Models trained with spectrogram features exhibit the weakest performance, with a Dev EER of 47.86% and Eval EER of 38.72%. This indicates that spectrograms, without additional learned enhancements, struggle to effectively separate genuine from fake audio in this task. Introducing Wav2Vec2 as the feature extractor significantly reduces the EER, even without pretraining, achieving a Dev EER of 38.24% and Eval EER of 40.64%. Despite this improvement, the lack of domain-specific pretraining likely limits its generalization capabilities.
Significant improvement is observed when the model is trained in two distinct phases. Initially, Wav2Vec2 undergoes self-supervised pertaining, as detailed in [18], on the ASVSpoof5 dataset, allowing it to learn meaningful representations of audio data without relying on labels. Subsequently, end-to-end training is performed using labeled data, fine-tuning both the feature extractor and the classification head for the specific task. This dual-phase training approach achieves a notable reduction in error rates, resulting in a Dev EER of 20.82% and an Eval EER of 26.36%, showcasing the combined benefits of unsupervised pretraining and supervised fine-tuning.
@@ 3.3 MaskedSpec and MaskedFeature Experiments
Squares 2.49 6.73
Bands 1.44 8.78
Singles 2.07 13.13
Gauss 1.38 5.22
@@
3.4 Comparison with Standard Augmentations
To further assess the effectiveness of our proposed augmentation strategies, we conducted an comprehensive series of experiments employing the most effective configurations of MaskedSpec and MaskedFeature, referred to as $MS_{B}$ (Bands) and $MF_{G}$ (Gaussian), respectively. In these experiments, we examined the performance of our augmentations alongside widely adopted techniques: RawBoost ($R$) and compression ($C$), the latter previously utilized in earlier subsections. These experiments were designed to assess not only the standalone benefit of our augmentations but also their capacity to enhance and complement existing methods.
Augmentation Configuration Dev EER (%) Eval EER (%)
None 3.20 8.34
$C$ 2.86 6.15
$R$ 5.56 6.11
$C$ + $R$ 3.51 6.14
$MS_{B}$ + $C$ 1.36 4.53
$MS_{B}$ + $R$ 2.29 4.82
$MS_{B}$ + $C$ + $R$ 2.86 4.99
$MF_{G}$ + $C$ 1.38 5.22
$MF_{G}$ + $R$ 3.26 5.21
$MF_{G}$ + $C$ + $R$ 1.88 4.94
Table 5: Equal Error Rates (EER) on the development and evaluation sets under various augmentation configurations. $MS_{B}$ = MaskedSpec (Bands), $MF_{G}$ = MaskedFeature (Gaussian) $C$ = Compression augmentation; $R$ = RawBoost. Best performance was achieved using $MS_{B}$ + $C$.
Table 5 presents model performance across various augmentation configurations. As a baseline, the model trained without any augmentation achieves a relatively high EER, particularly on the evaluation set (8.34%). Applying compression alone reduces the EER to 6.15%, while RawBoost achieves a comparable result (6.11%) but with a weaker Dev EER (5.56%). Interestingly, combining both compression and RawBoost does not yield further improvement, indicating potential redundancy or saturation effects between these augmentations.
In contrast, when our MaskedSpec ($MS_{B}$) or MaskedFeature ($MF_{G}$) augmentations are introduced, consistent gains are observed. Specifically, the combination of compression with $MS_{B}$ reduces Eval EER to 4.53%, while adding $MS_{B}$ to RawBoost results in a strong 4.82%. The same trend holds for $MF_{G}$, which lowers Eval EER to 5.22% and 5.21% when combined with compression and RawBoost, respectively. Notably, the combination $MF_{G}$ + $C$ + $R$ achieves better result at 4.94%, further highlighting the compatibility of our augmentation with existing techniques.
From a theoretical perspective, these improvements can be attributed to the nature of the distortions introduced. While RawBoost targets temporal variability by modifying amplitude and phase characteristics, and compression introduces global artifacts due to codec degradation, our augmentations inject structured masking along the frequency or feature axis. This masking enforces robustness to localized distortions and promotes better generalization across unseen manipulations. Thus, our methods operate on orthogonal principles relative to $C$ and $R$, providing complementary benefits when combined.
$MS_{B}$ + $C$ $MS_{B}$ + $R$ $MS_{B}$ + $C$ + $R$ $MF_{G}$ + $C$ $MF_{G}$ + $R$ $MF_{G}$ + $C$ + $R$ Eval EER (%)
3.74
4.02
3.82
3.98
3.92
4.01
Table 6: Fusion performance (Eval EER %) for models trained with different augmentation configurations. $MS_{B}$ and $MF_{G}$ denote MaskedSpec and MaskedFeature, respectively. $C$ = Compression augmentation, $R$ = Rawboost. Each row represents a distinct combination of fusion sources. The best performance achieved when fusing all six augmentation configurations.
In addition to evaluating individual augmentation effects, we explore the impact of combining models trained under different augmentation settings via late score-level fusion. Each fusion configuration in Table 6 is composed of the same models and training process, trained with the respective augmentations indicated in the columns. The goal is to assess whether diverse augmentation conditions lead to complementary decision patterns that enhance overall detection performance when combined.
As shown, the full fusion of all augmentations achieves the lowest EER (3.74%), indicating strong complementarity among the models. Removing any individual augmentation consistently degrades performance. This suggests that each component introduces unique perturbations that contribute valuable diversity in the learned feature space. The results highlight the benefit of leveraging augmentation diversity not only during training, but also in inference via ensemble-based strategies.
3.5 Different Feature Extractor Pretraining
Pretrained initialization of the feature extractor plays a significant role in shaping how speech signals are represented during downstream training, particularly in end-to-end architectures. Differences in dataset scale, language diversity, and domain alignment of the pretraining corpus can lead to varied representational biases, ultimately influencing performance. To examine these effects, we evaluate several feature extractor initializations using a unified training pipeline that incorporates our proposed augmentation strategies: MaskedSpec with Band masking ($MS_{B}$), MaskedFeature with Gaussian masking ($MF_{G}$), and compression augmentations.
Feature Extractor Pretraining Compressions $MS_{B}$ $MF_{G}$ Dev EER (%) Eval EER (%)
@@ 3.5 Different Feature Extractor Pretraining
ASVSpoof5+ 5.12 4.44
5.62 5.27
Table 7: Performance comparison across pretrained feature extractors and augmentation strategies. $MS_{B}$ = MaskedSpec (Bands), $MF_{G}$ = MaskedFeature (Gaussian), and $C$ = Compression augmentation. Best evaluation EER scores are highlighted in bold.
Table 7 summarizes the performance of each model configuration. As a baseline, we assess models initialized with XLS-R 53 and ASVSpoof5, both trained without any augmentation. XLS-R 53, pretrained on over 50k hours of multilingual speech, offers a broad and diverse representation space. In contrast, the ASVSpoof5 pretrained weights, although task-specific and focused on spoofed and bona fide speech, is limited in both scale and language diversity. This constraint limits its capacity to produce generalizable embeddings, resulting in significantly higher error rates. We then apply our augmentation strategy across all initialization settings. For XLS-R 53, incorporating $MS_{B}$ leads to a 26.4% reduction in Eval EER, while $MF_{G}$ results in a 15.1% decrease. The improvements are even more substantial for ASVSpoof5, with $MS_{B}$ reducing the Eval EER from 20.51% to 12.66% (38.3% improvement), and $MF_{G}$ from 20.51% to 14.17% (30.9% improvement).
The ASVSpoof5+ configuration introduces an additional insight, where compression augmentation applied during pretraining. This technique simulates real-world distortions and helps mitigate the limitations of low-resource datasets. With ASVSpoof5+, Eval EER is further reduced to 4.44% with $MS_{B}$ and 5.27% with $MF_{G}$, corresponding to relative improvements of 64.9% and 62.8% compared to the standard ASVSpoof5 configuration. This highlighting the impact of early exposure to distortion in enhancing feature learning.
To broaden our analysis, we also examine several publicly available large-scale models. XLS-R 128, trained on 436k hours across 128 languages, achieves the best overall results for Dev EER with 0.29% with $MF_{G}$, and one of the best Eval EERs (4.74%) with $MS_{B}$. These outcomes underscore the benefit of wide linguistic and acoustic coverage in self-supervised pretraining. WavLM, although monolingual, is trained on 94k hours of English speech and delivers competitive performance, suggesting that scale alone can provide strong generalization capabilities when paired with effective augmentations. In contrast, the Large-960h model, trained on 960 hours of LibriSpeech, underperforms relative to other models, illustrating the limitations of narrow, domain-specific pretraining.
The full set of results, including experiments with RawBoost and evaluation with minDCF, are presented in the Appendix. [^2] Please refer to Appendix A for detailed results.
3.6 Fusion of Models
Fusion Config ASVSpoof5+ XLS-R 53 XLS-R 128 WavLM Eval EER (%)
$MS_{B}$ $MF_{G}$ $MS_{B}$ $MF_{G}$ $MS_{B}$ $MF_{G}$ $MS_{B}$ $MF_{G}$
$C$ only 4.08
$C$ only 3.53
$C$ only 3.69
$C$ only 3.19
$C$, $R$, C+RC{+}R ✓* 2.92
$C$, $R$, C+RC{+}R ✓* 3.14
$C$, $R$, C+RC{+}R (full fusion) ✓* ✓* 2.71
Table 8: Fusion performance across models trained with different augmentation combinations and pretrained feature extractors. Augmentations used include compression ($C$), RawBoost ($R$), and their combination ($C+R$). * Models pretrained with RawBoost on ASVSpoof5+ were excluded due to degraded performance. The best fusion result is achieved when combining all models.
Table 8 presents the evaluation results of fusion across models trained with different augmentation configurations and feature extractor initializations. The fusion is performed by averaging output scores from independently trained models. This analysis examines whether combining models trained under complementary augmentation strategies leads to improved detection performance.
We first explore fusion within a closed condition setting, using only models trained with ASVSpoof5 dataset. When both masking strategies ($MS_{B}$ and $MF_{G}$) are included, the fused system achieves an Eval EER of 4.08%, surpassing the performance of individual models in isolation.
We then expand the fusion scope to include models trained with different feature extractor initializations (XLS-R 53, XLS-R 128, WavLM), trained with compression. Fusing models trained individually with either $MS_{B}$ or $MF_{G}$ produces notable improvements (Eval EER of 3.53% and 3.69%, respectively). However, combining both masking strategies across all extractors further improves performance, yielding an Eval EER of 3.19%.
Finally, we extend the fusion to encompass models trained with compression, RawBoost, and their combination. While RawBoost had degraded results under ASVSpoof5+ and is excluded in those cases (marked with *), its inclusion from other pretraining sources contributes valuable temporal perturbation patterns. Fusion of models trained only with $MS_{B}$ or only with $MF_{G}$ gives Eval EERs of 2.92% and 3.14%, respectively. The full combination of both masking types across all augmentation regimes achieves the lowest error rate of 2.81%.
3.7 Cross-Dataset Evaluation
To rigorously assess the generalization of our framework under domain shift, we conduct cross-dataset evaluations and benchmark against leading systems reported in [20]. While most prior systems were trained on ASVSpoof2019, our primary configuration was trained on ASVSspoof5 (2024) (Track 1). This comparison highlights the effect of different training sets when models are evaluated beyond their source domain.
System Training Data ASVSpoof19 Eval ASVSpoof21 LA ASVSpoof21 DF
[…]
Rawformer ASVSpoof2019 0.59 4.98 4.53
Bi-LSTM + MLP ASVSpoof2019 1.28 6.53 4.75
DARTS ASVSpoof2019 1.08 7.89
Ours ASVSpoof5 (2024) 4.74 12.58 7.94
[…]
Table 9: EER (%) comparison on ASVSpoof2019 evaluation and ASVSpoof2021 LA and DF evaluation sets. The training dataset used for each system is explicitly indicated. Dark gray indicates the best-performing system, and light gray denotes the second-best.
Table 9 shows that when trained on ASVSpoof5 (2024), our system demonstrates reasonable resilience but does not reach state-of-the-art performance under cross-dataset evaluation. While these results are acceptable, we sought to determine whether the limitation stems from the model itself or from dataset mismatch. To this end, we trained the system on the ASVSpoof2019 training set, which is considerably smaller than ASVSpoof5, and compared it against current state-of-the-art systems. In this setting, our model achieved second-best performance on both the ASVSpoof2019 evaluation and the ASVSpoof2021 DF tasks, with EERs of 0.18% and 2.92%, respectively. These results underscore the effectiveness of our augmentation framework even when trained on a substantially smaller dataset.
From these results, we conclude that: 1. The relatively weaker performance of ASVSpoof5-trained models is primarily due to dataset mismatch rather than inherent shortcomings of our framework. 2. Resilience to domain shift remains a central challenge for practical deployment. We believe that designing feature extractors capable of learning general and invariant representations is key to improving robustness and ensuring reliability in industrial systems.
4 Analysis of Model Performance
In this section, we present a detailed analysis of our best preformed systems using the evaluation methods provided by the ASVspoof5 challenge organizers. We examine aggregated trends in EERs across a range of spoofing attacks types and codecs conditions, allowing us to infer model behavior. By analyzing pooled EERs per condition, we could identify patterns of vulnerability and resilience, assess the effectiveness of our augmentation strategies, and highlight which model configurations offer the most consistent performance across diverse real-world scenarios.
4.1 Spoofing Attacks
@@
XLSR53, $MS_{B}$ + C 1.62 2.68 1.32 1.33 1.56 2.34 3.53 6.03 1.42 1.41 1.76 17.32 1.42 3.76 5.10 1.47
XLSR53, $MF_{G}$ + C 2.00 4.53 2.01 1.75 2.22 2.58 5.05 8.71 1.63 1.94 2.09 16.30 1.74 5.52 6.60 1.55
XLSR128, $MS_{B}$ + R 0.96 2.77 2.33 2.36 0.92 1.46 2.99 3.06 1.55 1.50 5.12 8.99 0.75 7.01 6.34 3.57
XLSR128, $MS_{B}$ + R + C 0.71 2.15 2.43 3.70 0.66 1.35 2.91 3.04 1.54 1.41 6.34 6.05 0.58 6.63 6.50 6.12
XLSR128, $MF_{G}$ + C 1.86 4.18 1.79 1.48 1.14 1.76 2.61 5.81 1.73 1.81 3.60 12.74 1.42 6.96 6.46 2.47
XLSR128, $MF_{G}$ + R + C 1.05 2.77 1.76 2.92 1.18 1.50 4.67 3.58 1.61 1.63 5.50 9.18 0.77 6.41 7.53 3.96
ASVSpoof5, $MS_{B}$ + C 1.69 0.84 6.20 3.42 2.29 3.10 1.63 4.55 3.83 1.45 1.25 16.14 1.08 1.19 3.10 1.62
ASVSpoof5, $MF_{G}$ + C 2.02 0.82 6.49 5.34 1.85 4.30 1.47 5.57 5.20 1.37 1.34 18.22 1.14 1.49 4.60 2.13
ASVSpoof5+ fusion 1.16 0.58 5.23 3.51 1.53 2.66 1.19 4.24 3.34 1.22 1.13 15.01 0.83 0.94 2.94 1.56
Full fusion 0.55 0.94 1.58 1.58 0.89 1.18 1.68 3.53 0.91 1.06 1.82 7.86 0.54 1.88 3.33 1.49
@@
Table 10: Pooled EER performance analysis results for a selected subset of top-performing models across 16 spoofing attack types (A17–A32). Bold values indicate the lowest EER for each codec among the single models, while light gray denotes the second-best.
The ASVSpoof5 evaluation set introduces a diverse range of unseen spoofing attacks, labeled A17–A32. Table 10 presents the pooled EERs for each attack across various model configurations. As observed, attacks such as A17, A21, A22, A25, A26, and A29 consistently result in low EERs across most models, suggesting that these spoofing types contain identifiable artifacts or exhibit acoustic patterns that are effectively addressed by the applied augmentations. Attacks A24, A28, A30, and A31 present more substantial challenges, as evidenced by elevated error rates in nearly all configurations. Among them, A28 stands out as the most difficult to detect, with numerous models exceeding 10% EER, indicating its high perceptual similarity to bona fide speech.
Model initialized with XLS-R 53 and trained using $MS_{B}$ + $C$ exhibit strong generalization, achieving low EERs on a diverse set of attacks, including A19, A20, A25, and A32. In particular, XLS-R 128 combined with $MS_{B}$ + $R$ + $C$ delivers leading performance across both moderate and high-difficulty spoofing attacks scenarios, including A24 and A28. This supports the value of leveraging extensive multilingual pretraining along with complementary augmentations for enhanced robustness.
Interestingly, models pretrained on the more constrained ASVSpoof5+ dataset outperform larger models on several spoofing types where general-purpose pretraining appears less effective. Specifically, the $MS_{B}$ + $C$ variant achieves the lowest EER on A27, A30, and A31, while $MF_{G}$ + $C$ outperforms others on A18, A23, and A26. These outcomes indicate that domain-specific pretraining, even at smaller scale, can offer competitive advantages when paired with task-relevant augmentations, likely due to improved alignment with the spectral and temporal characteristics of fake audio.
4.2 Codecs
Model None C01 C02 C03 C04 C05 C06 C07 C08 C09 C10 C11
XLSR128, $MS_{B}$ + C 0.59 2.17 2.14 3.62 12.13 0.72 1.18 16.29 5.43 3.61 5.04 1.06
XLSR128, $MS_{B}$ + R 1.44 2.62 2.43 3.60 8.59 1.67 2.12 11.03 4.37 3.80 5.20 1.70
XLSR128, $MS_{B}$ + R + C 1.10 2.54 3.08 3.35 9.17 1.52 1.77 12.51 4.13 4.02 5.60 1.21
XLSR128, $MF_{G}$ + R+ C 1.84 2.89 3.21 4.65 8.31 2.59 2.56 11.07 4.58 4.05 4.60 1.83
WAVLM, $MS_{B}$+R+C 1.00 3.66 2.11 4.12 9.27 1.25 2.48 12.77 4.55 5.53 6.17 1.63
ASVSpoof5, $MS_{B}$ + C 3.35 4.09 3.62 5.07 4.97 4.13 4.29 5.85 5.87 4.76 5.55 3.42
ASVSpoof5, $MF_{G}$ + C 3.83 4.83 4.44 5.98 6.05 5.19 4.92 7.09 7.24 6.33 6.65 3.82
AVSpoof5+ fusion 2.98 3.61 3.28 4.64 4.68 4.09 4.04 5.64 5.34 4.57 5.15 2.98
Best fusion 0.44 1.39 0.95 2.37 6.04 0.62 0.97 8.21 2.87 2.19 3.30 0.90
Table 11: Pooled EER performance analysis for a selected subset of top-performing models across 12 codec conditions (C01–C11) and the uncompressed baseline (“None”). Bold values indicate the lowest EER for each codec among the single models, while light gray denotes the second-best.